OpenStudy (anonymous):

A truck is rented at $40 per day plus a charge per mile of use. The truck traveled 15 miles in one day and the total charge was $115. Which equation can be used to find the charge x, in dollars, per mile traveled? 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115 40 + 15 = 115x

OpenStudy (luigi0210):

Do you know how to start?

OpenStudy (anonymous):

no

OpenStudy (luigi0210):

Well we know that the total for one day was $115, so that doesn't change. Also it says $40 per day, and the question states it was for one day so that stays the same. Leaving..

OpenStudy (luigi0210):

What are we left with?

OpenStudy (anonymous):

well i have 40x + 15 = 115 40 + 15x = 115 40x + 15x = 115 40 + 15 = 115x to choose from

OpenStudy (luigi0210):

Do you just want ansurs?

OpenStudy (anonymous):

well for this one kinda i normally wouldnt no tho but yes for this one.

OpenStudy (anonymous):

thank you. :/

OpenStudy (luigi0210):

Yup, btw Welcome to openstudy :)